Please anyone can help me with this probability question and please provide explanation. Thank you so much!

  1. (i) Consider two events A and B, with P(A) = 0.3, P(B) = 1. Compute P(A∩B), P(Ac ∩B), and P(A ∩ Bc) (where we denote by Ac = Ω \ A the complement of an event A).

    (ii) We now consider three independent events A, B, and C. Using the definition of inde- pendence, show that the two events Ac and Bc ∪ Cc are independent.
Answer #1

ACAB Sch. W Given P(A) = 0.3 , P(B) = 1 Clarly ACB So AnB = A @ P(ANB) = P(A) = 0.3 And OPAB) = P(B) – PCA) = 1 - 0.3 = 0.7 A